OpenStudy (anonymous):

plzzz i need your smart brains for this Solve for x. 6/x=x/12 the answer could be 36 2 √6 6 √2 72

OpenStudy (anonymous):

start with \(x^2=6\times 12=72\)

OpenStudy (anonymous):

so the ans is 72?

OpenStudy (anonymous):

thanks alot dude :)

OpenStudy (anonymous):

oh no it says \(x^2=72\) so \(x=\sqrt{72}\)

OpenStudy (anonymous):

oh lol okay

OpenStudy (anonymous):

6/x=x/12 12x6=x^2 sqrt(3x2x2x2x3)=x^2 3x2xsqrt(2)=x \[6\sqrt{2}\]=x

OpenStudy (anonymous):

or \[\sqrt{72}=\sqrt{36\times 2}=6\sqrt{2}\]

OpenStudy (anonymous):

oh thank you guys alot !!!!!

OpenStudy (anonymous):

answer is c

OpenStudy (anonymous):

thanks